OpenStudy (abbycross167):

Can someone please show me the steps to answer this question? A line with slope 1/3 passes through the point (3, 6). Which point also lies on this line? a. (6, 3) b. (7, 6) c. (-3, -3) d. (-6, 3)

OpenStudy (abbycross167):

@AloneS @TheProfessor35 can one of y'all please help me?

OpenStudy (theprofessor35):

Put this into point slope form? Do you know the formula?

OpenStudy (theprofessor35):

y - y1 = m(x - x1)

OpenStudy (theprofessor35):

Lets plug the stuff into the formula now.

OpenStudy (theprofessor35):

Could you do that? Hint: m = slope (x1,y1)

OpenStudy (abbycross167):

y- 6 = 1/3(x- 3)

OpenStudy (theprofessor35):

Nice job! Now lets solve.

OpenStudy (theprofessor35):

y- 6 = 1/3(x- 3) y - 6 = 1/3x - 1 y = 1/3x + 5

OpenStudy (theprofessor35):

Oh my....I lost track....I don't know where to end this.......... Sorry :(

OpenStudy (abbycross167):

Wait what?

OpenStudy (abbycross167):

@TheProfessor35 what do you mean?

OpenStudy (theprofessor35):

I don't think I know how to finish it..... :(

OpenStudy (abbycross167):

oh... its ok, do you know anyone that could help me?

OpenStudy (abbycross167):

@mustbethestars do you think you could help me?

OpenStudy (abbycross167):

@Ineedhelplz could you try to help me?

OpenStudy (ineedhelplz):

Is there a base?

OpenStudy (ineedhelplz):

btw I got d

OpenStudy (ineedhelplz):

I graphed it

OpenStudy (abbycross167):

could you please show me how you got d? @Ineedhelplz

OpenStudy (ineedhelplz):

I graphed it!

OpenStudy (mustbethestars):

The graph looks off to me.

OpenStudy (ineedhelplz):

The red dot is 3,6 and the orange one is -6,3

OpenStudy (ineedhelplz):

@mustbethestars How?

OpenStudy (abbycross167):

Yes sir/ma'am but could you show me the step to get -6,3?

OpenStudy (ineedhelplz):

I plugged in the base and the slope y=mx+b

OpenStudy (abbycross167):

so y=1/3 3+6?

OpenStudy (ineedhelplz):

I got 5 for the base

OpenStudy (ineedhelplz):

y=1/3(.3(repeating))x+5

OpenStudy (ineedhelplz):

y=1/3x+5

OpenStudy (abbycross167):

Wait I'm confused

OpenStudy (abbycross167):

how did you get 5?

OpenStudy (abbycross167):

@Ineedhelplz

OpenStudy (abbycross167):

@mustbethestars do you know how to do this?

OpenStudy (mustbethestars):

I'm sorry I have no idea

OpenStudy (abbycross167):

oh its ok.... do you know anyone who could help me?